# A Literature Review of Literature Reviews in Pattern Analysis and Machine Intelligence

### v4

@@ 4.2.1 Reference Quality Measurement (RQM)
The reference quality of a literature review is a multifaceted concept. It usually involves several direct factors such as credibility, relevance, breadth, and depth, etc. Quantitatively evaluating these direct elements poses significant challenges. On the one hand, the relevance of scientific literature is difficult to be precisely defined. Relying on co-citation analysis [92] or the similarity of paper embeddings both suffer from conceptual limitations. On the other hand, practical limitations are evident. For instance, the concept of breadth, also known as coverage, within a literature review, can theoretically be quantified as the ratio of the number of references to the total number of relevant references. However, accurately determining this ratio is challenging. It is difficult to ascertain the complete number of relevant references either by keyword search or by citation network.
Due to the challenges of quantifying the reference quality using the above-mentioned direct factors, we consider an indirect quality indication of each reference with the assistance of the $TNCSI$ presented in Sec 4.1.1. Here, the $TNCSI$ can be regarded as a reference quality indicator based on user voting, which is a statistical result of numerous researchers' comprehensive analyses of those direct factors of a certain paper.
Timeliness also matters. A current and up-to-date literature review ensures that the most recent advancements, developments, and perspectives in a particular field are taken into consideration. This temporal relevance enhances the accuracy and effectiveness of research outcomes, as it reflects the current state of knowledge and understanding. By focusing on the currency of the references, we can gauge the extent to which the literature review incorporates the latest research and developments in the field.
@@
To simultaneously consider the quality and timeliness of cited references, we propose modifying the Gompertz function to model the reference quality (see in Eq. (11)). The Gompertz function is characterized by its sigmoidal, which indicates a slow growth rate at the start and end of a time period, with a more rapid growth in the middle phase. This pattern is often observed in natural phenomena, such as species dynamics [16], tumor growth [99], etc.
$RQM=1-e^{-\beta\cdot e^{-(1-ARQ)\cdot S_{mp}}}$ (11)
where $\beta$ is the shift parameter, $ARQ$ stands for average reference quality, and $S_{mp}$ represents the median semester count of the reference age [94], defined as the period spanning from the publication dates of the cited references to the issuance of the review.
@@
The calculation procedures of $ARQ$ are as follows: The first step is the extraction of the cited reference list. For most publications, their reference lists could be provided by Semantic Scholar API. For a small number of reviews, the reference list provided by Semantic Scholar may contain errors. In this case, although there are powerful computer vision-based algorithms [14, 22] available for extracting the reference list within PDFs, our requirements are relatively simple and can be effectively met by relying on the heuristic algorithms or ChatGPT. More specifically, for literature review with a relatively fixed citation format, we can use the PDFMiner [79] to read text from PDF files and use heuristic rules to match citations. Alternatively, the text can be analyzed using ChatGPT to extract in-text citations. The second step is similar to the approach presented in Sec. 4.1.1, where the ChatGPT and a well-designed prompt (as presented in Fig. 5) are utilized to obtain the topic keyword of the review. Next, we calculate the $TNCSI$ for each reference in the list. To conserve computational resources, we avoid using the ChatGPT to generate keywords for each reference. Instead, the $TNCSI$ of all cited literature is calculated using a sharing topic keyword. Finally, the coverage can be further calculated in Eq. (12):
$ARQ=\frac{\sum_{i=1}^{N_{R}}TNCSI(Ref_{i})}{N_{R}}$ (12)
@@
The shift parameter $\beta$ can be set empirically or obtained statistically. For statistical estimation, we first examine the distribution of $S_{mp}$ and $ARQ$ of the RiPAMI database. The results indicate that (1) the $S_{mp}$ of over 50% of the reviews falls within the [5,10] interval, where we denote the lower boundary as $l_{s}$ and the upper boundary as $r_{s}$; (2) the average of $ARQ$ is approximately 0.6, denoted as $\overline{ARQ}$. Then, the problem of asserting for $\beta$ is reconceptualized as an optimization problem. As shown in Eq. (13), the objective here is to identify the value of $\beta$ that maximizes the integral of the absolute value of $RQM^{\prime}(S_{mp};\beta,\overline{ARQ})$ over the range $l_{s}$ to $r_{s}$, subject to the constraints of $\overline{ARQ}=0.6$. Such an approach would endow $RQM$ with a more discriminative nature. It should be noted that different fields may result in distinct values of $\beta$. For this study, the $\beta$ has been established as 5.
$\beta_{\text{opt}}=\arg\max_{\beta}\left(RQM^{\prime}(\overline{S_{mp}};\beta,\overline{ARQ})\right)$ (13)
The range of $RQM$ extends from 0 to 1, where values closer to 1 signify a higher quality of the referenced literature. As illustrated in Fig. 6 (a), when the $ARQ$ of a paper remains constant, an increase in the variable $S_{m}p$ will lead to a decrease in the $RQM$ value. Conversely, when $S_{m}p$ remains constant, a higher $ARQ$ will elevate the $RQM$ value.
4.2.2 Review Update Index (RUI)
@@
$CDR=\frac{N_{pc}\cdot R_{k}}{R_{k}\cdot N_{mp}}=\frac{N_{pc}}{N_{mp}}$ (14)
@@
Finally, the $RUI$ could be obtained by weighted summation of $CDR$ and $RAD$:
$RUI=p\cdot CDR+q\cdot RAD(M_{pc})$ (16)
where $p$ and $q$ are set to 10 and 5 in this paper, respectively. A sculptural visualization of the $RUI$'s contours is crystallized in Fig. 6 (b).
Figure 6: 3D visualization of the proposed quality indicators. Panels (a) and (b) respectively depict the value landscapes of $RQM$ and $RUI$, shaped by two independent variables.
4.3 Quantitative Evaluations of Literature Reviews
We have carefully selected numerous reviews within the field of PAMI to ensure a representative sample for our evaluation. By employing the proposed quantitative indicators, we conduct a thorough assessment of these scholarly surveys, aiming to gauge their quality and potential influence within the field.
In Fig. 7, we present a set of scatter plots depicting the references' quality of five randomly selected literature reviews published in two representative journals between 2015 and 2018. To prevent conflicts of interest, we use Journal A to represent a journal with an impact factor greater than 20. Journal B represents a journal with an impact factor below 5. The reference quality visualizations of journal A and journal B are presented in Fig. 7 (a), (b), respectively. The horizontal axis represents the reference age, and the vertical axis corresponds to the proposed $aTNCSI$. The color and size of the scatter points indicate the $IEI$ of the references. A positive $IEI$ value signifies a gradually increasing citation trend, resulting in warm-colored scatter points. The larger the size of the scatter point, the greater the value of $IEI$. Conversely, when the $IEI$ value is negative, the scatter points are cool-colored, and the size decreases as the value decreases. A closer examination of Fig. 7 reveals distinct distribution patterns among various journals depicted in their respective scatter plots. Journal with higher impact factors tends to exhibit clustered dots in the upper left corner, indicating a trend toward referencing more recent and influential sources. Conversely, a journal with lower impact factors is prone to display more dispersed dots, with a notable distribution in the lower half of the graph. We argue that this discrepancy stems from authors' tendency to cite recent papers in their literature reviews. Given that these relatively recent papers only had limited received citations, their citation count might not adequately reflect the potential value of the literature. Therefore, those seasoned researchers submitting reviews to top-tier journals often render more precise judgments about the prospective significance of a certain paper, as evidenced in the figure.
Detailed numerical results can be found in Tab. IV, where ``Cites'' represents the number of received citations, and ``Refs'' denotes the number of references. As can be seen from the table, literature reviews [7, 18] of different topics that receive approximately the same number of citations exhibit notable discrepancies in their associated $aTNCSI$. This divergence primarily stems from the varying levels of research interest across their respective topics. Surveys [101] in some emerging fields reveal a significant increase in their $IEI$, as expected. The majority of leading journals demonstrate high $RQM$ values, which reflect the effectiveness of high-standard peer review. In addition, even if the topics covered by the reviews [121, 119] are similar and their publication dates are relatively close, the proposed $RUI$ still provide some indication of the extent to which they need to be updated. Note that, given the time span of less than six months, the calculation of $IEI$ for certain papers is unavailable (denoted with ``-'').
Figure 7: Visualization of references quality of randomly selected reviews published in different journals. Panel (a) displays the citation quality distribution for a journal with an IF Score above 20, while Panel (b) shows the same for a journal with an IF Score below 5. This suggests that reviews published in journals with higher impact factors tend to reference sources of superior quality and greater timeliness.
@@ 4.3 Quantitative Evaluations of Literature Reviews
TABLE IV: Comparison of various reviews with the proposed indicators and metrics.
@@
5 Human-authored vs. AI-generated
5.1 Overview of AI-generated Literature Reviews
Traditionally, literature reviews have been manually conducted by researchers who analyze and synthesize scholarly sources to provide an overview of the current state of knowledge in a specific field. Recently, with the advancement of AI technologies, there's been a growing interest in leveraging artificial intelligence techniques, especially large language models, to automate or assist in the generation of the literature review. Typically, users are simply required to indicate their area of research interest, and the system will then automatically generate a literature review.
The automated creation of a literature review is a multidisciplinary endeavor that integrates knowledge from various fields. It relies not only on artificial intelligence technologies but also requires the merging of knowledge from other fields such as data science, bibliometrics, database engineering, etc. These components are instrumental in extracting, storing, and synthesizing relevant information from vast amounts of literature.
Early attempts in AI-generated literature reviews involve training language models, e.g. LLama [97], on a large corpus of academic papers, research articles, and other scholarly content. These models can then be used to generate coherent and contextually relevant text based on prompts or queries related to a specific research topic. However, given that a fully trained or fine-tuned language model has no access to the latest scholarly advances, these models would not generate a literature review that includes the latest scholarly materials. Thus, most existing advanced AI-generated literature review systems contain three main steps: knowledge retrieval, synthesis, and report.
Though researchers are expecting to save significant time and effort in conducting comprehensive surveys by embracing AI-generated literature reviews, numerous concerns about ethical issues and fabrications in AI-generated academic content have been raised. A critical appraisal by Zybaczynska [126] highlights that current AI systems typically fall short in providing substantial, accurate information and critical discernment. Elali et al. [28] critically examines the profound challenges brought by the fabrication and falsification of AI-generated research, underscoring its significant impact on the scientific community. Furthermore, numerous academic publishers are cautious about content created by AI. For instance, journals such as ``Nature'' and ``Science'' prohibit listing ChatGPT or other automated tools as authors on papers published in their issues. Nevertheless, we argue that the automated generation of literature reviews not intended for publication remains meaningful. It may assist researchers in staying rapidly abreast of the latest advancements in thriving fields.
5.2 Current State of AI-generated Literature Review
Most of the existing systems are only capable of generating literature reviews in plain text form and with a relatively fixed layout. A popular Github repository ChatPaper [43] retrieves papers based on a user-defined keyword with the use of the Google Scholar crawler. After a cosine similarity-based filtering of the retrieved papers, the selected papers are processed sequentially using ChatGPT. The output of Chatpaper is brief and concise, which usually contains a narrative description of the selected papers. Paper Digest [77] is an AI-based online platform that provides the service of automated review generation. Users can generate a literature review by setting specific keywords and further refining the cited source through constraints on publication dates. Similar to ChatPaper, the output of Paper Digest is also a narrative plain text description. Not all systems for generating literature reviews are fully automated. For example, Jenni AI [4] requires user interaction through ``AI Command'' during the review creation process, enabling the generation of highly customized literature reviews. Additionally, there are numerous online platforms [86, 15] and plugins available in the GPT Store that offer automated literature review generation services. Given the constraints on manuscript length, detailed introductions to these systems will not be provided.
Despite previous discussions indicating that certain publishers refuse to accept non-human authors, scholars have still successfully published AI-generated literature reviews in some of academic journals. Aydın et al. investigate how well can LLMs perform in the generation of the literature review. They employed well-known LLMs including ChatGPT and Google Bard to generate reviews on digital twin in healthcare [9] and meta-universes [8]. In their attempts, the Google Bard (or the ChatGPT) was first adopted to paraphrase the abstracts of papers within the last 3 years. Then, they designed several question prompts (such as ``What is digital twin'') to query the LLMs and rearranged the response in a formatted style.
While promising progress has been achieved in AI-generated literature reviews, there remains considerable room for improvement. The pursuit of automated literature reviews that rival the quality of those written by humans is an ongoing area of research, underscoring the need for continued exploration and development in this field. For example, most existing literature review generation systems are not capable of extracting information from tables or figures. As a result, tables and figures will not be included in the generated reviews. More details about the differences between AI and human-authored reviews will be presented in the next subsection.
5.3 Case Study: Comparing Human and AI in Conducting Literature Reviews
We crafted a case study aimed at elucidating the distinctive approaches employed by human researchers and their AI counterparts. To prevent potential information leakage and ensure a fair comparison due to the possibility of ChatGPT having been exposed to relevant reviews during training, we selected a newly emerging field, i.e. visual prompt engineering, which developed after September 2021 (the cutoff date for ChatGPT's training), as the focus of our investigation. Considering that literature reviews generated by AI lack academic influence, we only calculate $RQM$ (as introduced in Sec. 4.2) of them.
The data presented in Tab V reveals that, at present, there is still a gap in knowledge retrieval, synthesis, and reporting steps where AI systems lag behind human performance. Despite the high $RQM$ achieved by the ``PaperDigest'' and ``askyourpdf'', the overall quality of the generated reviews still falls short of the human-authored counterpart, especially given its reliance on only less than 10 references. Additionally, to our knowledge, few automated review systems can seamlessly integrate visual elements such as figures and tables. This landscape of predominantly plain text-based AI-generated reviews invites intriguing possibilities for future exploration and enhancement in the field.
It is noteworthy that despite the higher quality of the human-authored review [100], the timeline from draft to journal acceptance, as indicated by its preprint publication and acceptance dates, extends over six months. In contrast, AI systems can generate reviews within seconds. This delay implies that in rapidly advancing research areas, some of the newest findings might be overlooked by human authors.
Review Reference RQM↑ Automation Level Visual Elements SALSA Analysis [31]
[100] by human authors 160 0.96 Manullay
A typically narrative, method clustering-based review aims to offer valuable insights in visual prompt learning. The selection criteria for references are not specified, but it's clear that each reference has been thoroughly appraised.
Review by Jenni [4] - - Semi-automated $\times$
Automated searching and appraising references is not supported. The generation process relies on user interaction, and only narrative description is provided.
Review by ChatPaper [43] 15 0 Automated $\times$
Retrieving relevant literature from arXiv based on multiple LLM-generated keywords without appraisal step. Each section contains plain descriptions of the related reference.
Review by PaperDigest [77] 5 0.92 Automated $\times$
Searching papers with the user-specified keyword. It seems to appraise the quality of references with private criteria. The generated content is more like a concise summary.
Review by askyourpdf [15] 9 0.99 Automated $\times$
No official explanations are found for how to retrieve references and appraise the quality of the literature. The generated review includes a brief analysis and description of the related concept, current state of development, and gaps.
TABLE V: Comparisons between human-authored and AI-generated literature reviews.
6 Challenges and the Future of the Literature Review
6.1 Challenges
High Rate of obsolescence Knowledge in scientific research continually evolves and updates. Literature reviews require significant time and effort to compile. Due to the slower pace of review crafting, they may not reflect the latest research advancements in a timely manner, especially when these reviews are conducted by human authors and in need of a peer-review process.
Information Overload While our proposed indicators mitigate the problem of information overload to a certain degree, several challenges still await resolution. With the continuous growth of scientific research, the volume of literature is rapidly expanding. Collecting and screening a large number of publications within those blooming fields would cause inevitable incomplete searches and other similar issues. Both human and AI systems have to explore how to retrieve relevant literature more comprehensively and effectively.
Bias Literature reviews conducted in a specific language or region may unintentionally omit relevant studies published in other languages or regions. This language and geographical bias can limit the global perspective and generalizability of the review's findings. Such a bias appears not only in literature reviews written by humans but also in those generated by AI systems (even featuring multilingual capabilities).
6.2 Future
AI Empowerment Dynamic Literature Review In the rapidly advancing landscape of scientific research, there is a growing trend toward dynamic reviews generated in real time. While debates persist regarding the adherence of AI-generated content to academic ethics and its suitability for formal publication, the generated reviews for non-publishing purposes remain widely embraced. Given the labor- and time-intensive nature of conducting literature reviews, the exploration of AI technology for the automated generation of dynamic literature reviews stands as an area warranting further investigation.
Automated Literature Appraisal Literature appraisal is important for both human authors and AI systems. Although four quantitative indicators are proposed in this paper, this still makes it difficult to fully capture the academic impact and the quality of a certain review paper. In the future, with more powerful large multimodal models (LMMs) [50, 3], it will be possible to extract valuable information from non-textual modal content (e.g., images and tables) for automated literature appraisal.
Open Science and Advanced Search Engines The open science movement endorses the sharing and accessibility of literature data, thereby providing a greater number of resources for analysis and synthesis. It is recommended that the percentage of open-access papers should be further increased in the era of e-publishing. Furthermore, search engines that rely on semantic similarity, recommend systems, or co-citation networks rather than keyword matching are also encouraged. Such engines will primarily benefit both researchers and AI systems by enabling them to retrieve more relevant papers, thus enhancing the quality of any publications in all of the research fields.
7 Conclusion
This paper has provided a fresh perspective on the wealth of literature reviews in PAMI, introducing a systematic approach to categorize and evaluate them. In total, it presents four large language models-empowered quantitative evaluation indicators, a subjective evaluation that includes a typology for literature reviews, a comparison between human-authored and AI-generated reviews, and a meta-data database named RiPAMI, accompanied by a dataset of review topic key phrases.
The proposed evaluation indicators offer an innovative alternative to traditional bibliometric analysis, enabling a cost-efficient, field-normalized, and real-time assessment of literature reviews' quality and impact. These bibliometric indicators not only furnish clear numerical hints for human researchers but could also offer substantial assistance in the appraisal processes of AI-generated review systems.
Subject evaluations for reviews in PAMI are also provided. Numerous representative reviews of popular research topics are investigated to offer readers a brief overview of these exemplary reviews. In addition, a typology for reviews is introduced. It category the reviews into three major types which are method clustering-based, challenge-oriented, and hybrid literature review. Such typology enhances the understanding of how literature reviews are organized and synthesized.
The study highlights the differences between human-authored and AI-generated literature reviews, pinpointing significant gaps in retrieval, synthesis, and reporting capabilities of AI-generated reviews compared to those crafted by humans. The insights gained from this comparison not only help to understand the current state of AI-generated reviews in PAMI but also suggest how they might evolve with advancing technology.
By constructing the RiPAMI database, we analyze approximately 3,000 review samples in the field of PAMI, which provides statistical support for the proposed quantitative indicators. Furthermore, a dataset of topic keywords is manually annotated to validate the effectiveness of various prompts, enabling the selection of the optimal prompt for more accurate data retrieval.
To further support and expand upon this work, we release the code framework that encompasses functionalities including metadata retrieval, indicator calculation, and data analysis, etc. While it is designed for analysis of literature reviews in PAMI, the framework is universally adaptable, catering to the expansive needs of researchers from disparate fields.
Appendix A Visualization of the Proposed Impact Indicators.
To provide readers with a clear understanding of the proposed quality metrics, We render the graphical representations of $TNCSI$ and $IEI$ in Fig. 8 and Fig. 9, respectively.
As can be seen in Fig. 8, TNCSI equals the area under the probability density function curve, which is fitted with the use of maximum likelihood estimation. In Fig. 9, the horizontal axis labeled 0 to 5 inversely denotes the months prior to the current month, with 0 representing 6 months ago and 5 denoting the previous month. The corresponding $IEI$ is approximately $-0.76$, which indicates a slightly decreasing citation trend of the investigated review over the last six months.

### v5

@@ A.3 RQM
The quality of references in a literature review is a multifaceted concept involving various aspects, such as credibility, relevance, breadth, and depth. Quantitatively assessing these elements is challenging. To begin with, precisely defining the relevance of scientific literature is difficult; both co-citation analysis [111] and similarity in paper embeddings have inherent conceptual limitations. For instance, while the concept of breadth in a literature review can theoretically be measured as the ratio of cited references to the total number of relevant references, accurately calculating this ratio is challenging. Identifying all relevant references through keyword searches or citation networks remains elusive.
Given the challenges in quantifying reference quality through the direct factors mentioned, we instead consider an indirect indication of reference quality using $TNCSI$. Here, $TNCSI$ serves as a quality indicator based on collective user assessments, effectively capturing a statistical summary of numerous researchers’ evaluations of these direct factors for a given paper. Timeliness also matters. An up-to-date literature review ensures that the latest advancements, developments, and perspectives within a field are incorporated. This temporal relevance enhances the accuracy and effectiveness of research outcomes by reflecting the current state of knowledge. By emphasizing the currency of references, we can assess how thoroughly the literature review integrates the most recent research and developments in the field.
@@
To account for both the quality and timeliness of cited references, we propose modifying the Gompertz function to model the reference quality (see in Eq. (3)). The Gompertz function is characterized by its sigmoidal, which indicates a slow growth rate at the start and end of a time period, with more rapid growth in the middle phase. This pattern is often observed in natural phenomena, such as species dynamics [16], tumor growth [118], etc.
@@
The value of $RQM$ is determined by the average reference quality (ARQ), the shift parameter $\beta$, and the median age of references $S_{mp}$. The calculation procedures of $ARQ$ are as follows: The first step is the extraction of the cited reference list. For most publications, their reference lists could be provided by Semantic Scholar API. For a small number of reviews, the reference list provided by Semantic Scholar may contain errors. In this case, although there are powerful computer vision-based algorithms [13, 25] available for extracting the reference list within PDFs, our requirements are relatively simple and can be effectively met by relying on the heuristic algorithms or ChatGPT. More specifically, for literature review with a relatively fixed citation format, we can use the PDFMiner [95] to read text from PDF files and use heuristic rules to match citations. Alternatively, the text can be analyzed using ChatGPT to extract in-text citations. The second step is similar to the approach presented in Sec. 2.4.1, where the ChatGPT and a well-designed prompt (as presented in Fig. 9) are utilized to obtain the topic keyword of the review. Next, we calculate the $TNCSI$ for each reference in the list. To conserve computational resources, we avoid using the ChatGPT to generate keywords for each reference. Instead, the $TNCSI$ of all cited literature is calculated using a sharing topic keyword. Finally, the coverage can be further calculated in Eq. (14):
$ARQ=\frac{\sum_{i=1}^{N_{R}}TNCSI(Ref_{i})}{N_{R}},$ (14)
@@
The shift parameter $\beta$ can be set empirically or obtained statistically. For statistical estimation, we first examine the distribution of $S_{mp}$ and $ARQ$ of the RiPAMI database. The results indicate that (1) the $S_{mp}$ of over 50% of the reviews falls within the [5,10] interval, where we denote the lower boundary as $l_{s}$ and the upper boundary as $r_{s}$; (2) the average of $ARQ$ is approximately 0.6, denoted as $\overline{ARQ}$. Then, the problem of asserting for $\beta$ is reconceptualized as an optimization problem. As shown in Eq. (15), the objective here is to identify the value of $\beta$ that maximizes the integral of the absolute value of $RQM^{\prime}(S_{mp};\beta,\overline{ARQ})$ over the range $l_{s}$ to $r_{s}$, subject to the constraints of $\overline{ARQ}=0.6$. Such an approach would endow $RQM$ with a more discriminative nature. It should be noted that different fields may result in distinct values of $\beta$. For this study, the $\beta$ has been set as 5 for all fields.
[…]
The range of $RQM$ extends from 0 to 1, where values closer to 1 signify a higher quality of the referenced literature. As illustrated in Fig. 12, when the $ARQ$ of a paper remains constant, an increase in the variable $S_{mp}$ will lead to a decrease in the $RQM$ value. Conversely, when $S_{mp}$ remains constant, a higher $ARQ$ will elevate the $RQM$ value.
Figure 12: 3D visualization of the proposed $RQM$
A.4 RUI
@@
$CDR=\frac{N_{pc}\cdot R_{k}}{R_{k}\cdot N_{mp}}=\frac{N_{pc}}{N_{mp}},$ (16)
@@
Finally, the $RUI$ could be obtained by weighted summation of $CDR$ and $RAD$, as shown in Eq. 4. A sculptural visualization of the $RUI$’s contours is crystallized in Fig. 13.
Figure 13: 3D visualization of the proposed $RUI$
A.5 More Examples of Indicator Usage
Tab. XII provides more examples of how the four proposed indicators further assist researchers in refining their review selections. Compared to relying solely on titles, citation counts, and publication dates, the proposed metrics provide more information from various perspectives to assist researchers in review selections.
A.6 Ensuring Fair Use and Avoiding Metric Misinterpretation
No metric is perfect. Researchers should recognize the limitations of metrics and avoid their misuse or misinterpretation. All citation-based metrics, in particular, are subject to various biases, such as those influenced by the Matthew effect, which can inadvertently contribute to the persistence and dissemination of inaccuracies within academic literature [112].
For TNCSI. $TNCSI$ is primarily used to assess the cumulative impact of literature reviews and should not be used to evaluate the quality of literature reviews, especially those published in different years.
For IEI. $IEI$ solely represents citation trend changes over a specific past period and should not be extrapolated to predict future citation trends solely.
For RQM. As mentioned earlier, due to practical and conceptual constraints, $RQM$ considers only the average impact and timeliness of references, without accounting for the relevance of the references to the investigated topic. Consequently, there is a potential for manipulating this metric by citing influential references unrelated to the topic to inflate the $RQM$ score.
For RUI. As its name suggests, the $RUI$ is a metric used to assess the extent to which a review requires updating. It is not designed to quantify the degree of obsolescence of the perspectives or conclusions within the review.
Additionally, it is important to note that all the metrics are designed for article-level analysis of reviews and should not be used to infer the development of an entire field based on the metrics of a single article.
A.7 Alignment with the Leiden Manifesto
Principle $TNCSI$ $IEI$ $RQM$ $RUI$
Quantitative evaluation should support qualitative, expert assessment.
Measure performance against the research missions of the institution, group or researcher. $\times$ $\times$ $\times$ $\times$
Protect excellence in locally relevant research. N/A N/A N/A N/A
Keep data collection and analytical processes open, transparent and simple.
Allow those evaluated to verify data and analysis.
Account for variation by field in publication and citation practices.
Base assessment of individual researchers on a qualitative judgement of their portfolio. N/A N/A N/A N/A
Avoid misplaced concreteness and false precision.
Recognize the systemic effects of assessment and indicators.
Scrutinize indicators regularly and update them. N/A N/A N/A N/A
TABLE XI: Alignment of Leiden Manifesto Principles with Proposed Metrics: “✓” indicates alignment with the principle, “$\times$” indicates non-alignment, and “N/A” indicates principles not applicable.
Table XI shows the alignment of our proposed four metrics with the specific guidelines of the Leiden Manifesto. The metrics—$TNCSI$, $IEI$, $RQM$, and $RUI$—demonstrate strong adherence to most principles outlined in the manifesto. Specifically, all four metrics align with principles emphasizing the integration of quantitative evaluation with qualitative expert assessment, transparency in data collection and analysis, verification of data and analysis by those evaluated, accounting for field-specific variations, and avoiding misplaced concreteness and false precision.
The metrics do not align with the principle of measuring performance against the research missions of the institution, group, or researcher, as indicated by the “$\times$” in the corresponding row. We would like to point out that this reflects a deliberate trade-off made to balance engineering complexity. While more fine-grained measurements of performance would indeed enhance the objectivity and fairness of the metrics, they would inevitably increase the difficulty of data acquisition. In the future, with advancements in academic search engines and breakthroughs in natural language understanding, this issue is expected to be effectively addressed.
@@ A.7 Alignment with the Leiden Manifesto
TABLE XII: Comparison of various reviews with the proposed indicators and metrics.
@@
Appendix B Information Extraction Techniques
B.1 Word Counts
Word counting differs from character counting, where methods relying solely on regular expressions often yield imprecise results due to the complexity of word segmentation rules. To achieve precise word counting, we utilize the Python NLTK library for word tokenization and exclude all tokens representing non-alphabetic elements. As a result, the presented word count only includes alphabetic words, excluding numbers, contractions, and any special symbols.
B.2 Visual Elements Counts
[…]
Figure 14: LLM for figure and table caption extraction: This LLM RAG-based approach allows for the retrieval of figure and table captions at an acceptable cost.
B.3 Review Feature Extraction
This paper performs a statistical analysis of six features across more than 3,000 review samples, providing insights into factors such as compliance with PRISMA standards and the presence of application sections. Similar to visual element counting, these features are identified using LLM RAG-based information extraction methods.
Our proposed method for extracting review features is depicted in Fig. 15. The process begins with document analysis and recognition. Due to the inherent structure of PDF files, the machine-readable text order often diverges from the natural reading sequence. Additionally, elements such as captions and superscripts can introduce noise, disrupting semantic continuity during direct text extraction. To overcome these challenges, we propose utilizing the Nougat model [13] to process document images and convert them into structured text. This structured text enables efficient extraction of the Table of Contents (TOC) and content from individual sections.
Following the discussion in Sec. 3.1, we adopt a cost-effective approach during RAG by limiting the need for the LLM to process the entire text. In simple terms, specific sections of the article are selectively provided to the LLM to generate targeted responses. Alongside the title and abstract, the extracted introduction section and TOC are provided to guide the LLM in identifying whether the authors propose a new taxonomy or include a dedicated section for inclusion and exclusion criteria (indicating adherence to PRISMA guidelines) in the reviewed literature. To assess whether the authors discuss preliminaries, applications, or future challenges, we propose relying exclusively on TOC information, further minimizing costs. For benchmarking, the LLM is instructed to analyze extracted captions (illustrated in Fig. 14) to determine whether the authors have performed quantitative benchmarking of existing methods. All responses generated by the LLM are formatted into a structured 0/1 schema and stored in the RiPAMI database.
Figure 15: LLM for review features extraction: By utilizing OCR technology [13], we extract structured text and input the introduction section (Intro.), table of contents (TOC), and captions of visual elements into the LLM to derive review features.
Appendix C The Relationship between Review Features and TNCSI
In this study, we dedicate our efforts to analyzing and synthesizing six key types of content features in review articles within the PAMI field. It is important to emphasize that the presence or absence of these features in an article should not be interpreted as a direct measure of its overall quality or potential academic value. As shown in Fig. 16, there is no apparent correlation between the distribution of these six features and $TNCSI$.
Figure 16: Analysis of six review content features and their association with $TNCSI$: Gene expression heatmap illustrating the independence between six review content features (PRISMA, Application, Benchmark, Preliminary, Taxonomy, and Discussion) and $TNCSI$. Darker shades indicate a match with the feature, corresponding to a value of 1.
Appendix D Code Framework: PyBiblion
This study involved extensive coding and engineering practices during the processes of dataset construction and statistical analysis. To support this work, we developed an open-source code library, PyBiblion, designed for bibliometric and statistical analysis. As its name suggests, PyBiblion is a Python-based library designed for bibliometric and statistical analysis, drawing from the Greek root “biblion”, meaning “book” or “document” to emphasize its focus on scholarly literature.
PyBiblion offers several core advantages, one of which is lazy loading technology. Unlike most existing frameworks, PyBiblion implements lazy loading for paper information. This means that no network communication occurs during the initialization of an instance object; instead, data is only loaded when it is explicitly accessed. This mechanism helps to reduce the number of requests, alleviating server-side communication pressure and fostering a more equitable usage environment. The second advantage of PyBiblion lies in its user-friendliness. Extensive engineering optimizations have been implemented to integrate information retrieval and metric computation seamlessly. This design enables users to compute metrics such as the proposed $TNCSI$ and $RQM$ with a single line of code, offering a highly efficient and intuitive experience. In addition, PyBiblion integrates numerous practical features, including database support, multithreading execution, visualization tools, and statistical analysis capabilities, further enhancing its functionality and versatility.